OpenStudy (anonymous):

How do you solve this problem?? Any help is appreciated!! 4(x+1)^1/2 -5(x+1)^3/2 + (x+1)^5/2 =0

OpenStudy (anonymous):

did you google it?

OpenStudy (anonymous):

no..does typing it like that actually help find a way to solve it?

OpenStudy (anonymous):

well u could type in the properties of the problem and u might get help on how to solve it like a step by step whats the unit of the subject? amd firther more what kind of math is tht put in the answers to those questions and you may find what your looking for

OpenStudy (anonymous):

if (x+1)^1/2=y then {y^2 = x+1} 4y-5y^3+y^5=0 or y(y^4-5y^2+4)=0 y=0 is one answer, x=-1. y^2=z z^2-5z+4=0 delta=25-16=9 z1=(5-3)/2 =1 = y^2 = x+1 x=0 z2=(5+3)/2 =4 = y^2 = x+1 x=3

OpenStudy (anonymous):

Thank you soo much! Do you know how to solve this one too? \[\sqrt{x} - 3\sqrt[4]{x} = 4\]

OpenStudy (anonymous):

You are welcome! x^1/4 =y {x=y^4} domain of x: x>=0 {square} so y>=0 too. y^2-3y-4=0 delta=9+16=25 y=(3-5)/2 =-1. -1>=0 is false. We can't speak about x. y=(3+5)/2=4 4>=0 is true. so x=4^4=256

OpenStudy (anonymous):

Once again thank you!!! do you know how to do this one? \[\sqrt{6}x ^{2} + 2x - \sqrt{3/2} = 0\]

OpenStudy (anonymous):

Could you please type it again? you can show power for example x^1/2 or x^(-1/2)

OpenStudy (anonymous):

Sure! \[\sqrt{6}x^2 + 2x - \sqrt{3/2} = 0\] its square root of 6 and then x^2 plus 2x minus the square root of 3/2 =0

OpenStudy (anonymous):

it's ax^2 + bx + c =0 delta =4- 4(sqrt6)(-sqrt3/2) =4+4.sqrt(9)=4+12=16 x1=(-2-4)/2.sqrt(6) =-sqrt(6)/2 x2=(-2+4)/2.sqrt(6) = 1/sqrt(6) =sqrt(6)/6

OpenStudy (anonymous):

Is it OK?

OpenStudy (anonymous):

I dont really understand the third and fourth lines

OpenStudy (anonymous):

x1=[-b-sqrt(delta)]/2a x2={-b+sqrt(delta)]/2a

OpenStudy (anonymous):

it's a general formula too find roots of quadratic equation.

OpenStudy (anonymous):

If delta>0 then we have two unequal roots.. delta=0, x1=x2 . only one root. delta<o then we have no roots in real numbers.

OpenStudy (anonymous):

ohh ok!! thanks so much!you really dont know how much youve helped me haha.do you mind if i ask you one more?(:

OpenStudy (anonymous):

i become glad if i could help you:)

OpenStudy (anonymous):

\[(1/(x-1)) -( 2/x ^{2}) = 0\]

OpenStudy (anonymous):

at first x uneqal 1 and 0 is our domain.

OpenStudy (anonymous):

*(x-1)x^2 : x^2-2(x-1)=0 or x^2-2x+2=0

OpenStudy (anonymous):

delta=4-8 = -4<0 so I think we have no roots in real number.

OpenStudy (anonymous):

you have answer?

OpenStudy (anonymous):

no..ive been trying all these different methods and i keep on getting numbers like 2,1,-1 and i plug them back in and they dont come out to equal 0

OpenStudy (anonymous):

I rewrite your equation: please tell me is it true?

OpenStudy (anonymous):

1/(x-1) + 2/(x^2) = 0

OpenStudy (anonymous):

excuse me minus -

OpenStudy (anonymous):

correct. its like that on the page but to the left of the minus, its a fraction written like on top of each other and same thing with the 2/(over) x^2

OpenStudy (anonymous):

1 / [1/(x-1)] is it tru?

OpenStudy (anonymous):

and -2/[1/x^2]

OpenStudy (anonymous):

like 1/x-1 and -2/x^2 = 0 if that makes sense

OpenStudy (anonymous):

now I see correct form of your equivalent. thank you. I think my answer is true.

OpenStudy (anonymous):

so its an extraneous solution?

OpenStudy (anonymous):

it has imaginary roots. if our domain contain imaginary numbers: a+ib a,b belong to R. i = (-1)^1/2

OpenStudy (anonymous):

I think here our domain is R; Real numbers

OpenStudy (anonymous):

ok..i havent learned about that ha

OpenStudy (anonymous):

imaginaru nums? no problem. not impportant now. but solution in real numbers>

OpenStudy (anonymous):

if we have an equality, we can multiply both side in same number or term.

OpenStudy (anonymous):

I multiply both side in (x-1).x^2 to make a quadratic equation.

OpenStudy (anonymous):

before, I determine the domain. in fraction a/b , b could not equal to zero.

OpenStudy (anonymous):

then we solve our quadratic equation.

OpenStudy (anonymous):

thank youu!!:))